OpenStudy (anonymous):

Can someone help me please... Thanks A new cruise ship line has just launched 3 new ships; the Pacific Paradise, the Carribean Paradise and the Mediterranean Paradise. The Carribean Paradise has 28 more deluxe staterooms than the Pacific Paradise. The Mediterranean Paradise has 29 fewer deluxe staterooms than four times the number of deluxe staterooms on the Pacific Paradise. Find the number of deluxe staterooms for each of the ships if the total number of deluxe staterooms for the three ships is 881. The Pacific Paradise has______deluxe staterooms The Caribbean Paradise has____deluxe

OpenStudy (skullpatrol):

Step 1 Read the problem carefully. Decide what unknown numbers are asked for and what facts are known.

OpenStudy (anonymous):

Im still confuse with this problem.

OpenStudy (anonymous):

Can someone help me please... Thanks A new cruise ship line has just launched 3 new ships; the Pacific Paradise, the Carribean Paradise and the Mediterranean Paradise. The Carribean Paradise has 28 more deluxe staterooms than the Pacific Paradise. The Mediterranean Paradise has 29 fewer deluxe staterooms than four times the number of deluxe staterooms on the Pacific Paradise. Find the number of deluxe staterooms for each of the ships if the total number of deluxe staterooms for the three ships is 881. The Pacific Paradise has______deluxe staterooms The Caribbean Paradise has____deluxe The mediterranean Paradise has___deluxe staterooms.

OpenStudy (skullpatrol):

Decide what unknown numbers are asked for and what facts are known.

OpenStudy (skullpatrol):

Decide what unknown numbers are asked for. Find the number of deluxe staterooms for each of the ships

OpenStudy (anonymous):

28,29,881

OpenStudy (skullpatrol):

How?

OpenStudy (anonymous):

I still dont get this

OpenStudy (skullpatrol):

Let P = the number of rooms on the Pacific, since The Carribean Paradise has 28 more deluxe staterooms than the Pacific Paradise. Then C = 28 + P.

OpenStudy (anonymous):

ok

OpenStudy (skullpatrol):

The Mediterranean Paradise has 29 fewer deluxe staterooms than four times the number of deluxe staterooms on the Pacific Paradise. M =4P - 29

OpenStudy (anonymous):

k

OpenStudy (skullpatrol):

P, C, M P = the number of rooms on the Pacific C = 28 + P M =4P - 29

OpenStudy (anonymous):

ok

OpenStudy (skullpatrol):

Step 3 Reread the problem and write an equation that represents relations among the numbers in the problem.

OpenStudy (skullpatrol):

P + C + M = 881 if the total number of deluxe staterooms for the three ships is 881.

OpenStudy (anonymous):

3+28+4p-29=881

OpenStudy (skullpatrol):

Substitute into P + C + M = 881 the three: P = the number of rooms on the Pacific C = 28 + P M =4P - 29

OpenStudy (skullpatrol):

P + C + M = 881 P +(28 + P) + (4P -29) = 881

OpenStudy (anonymous):

so, p+(28+p)+(4p-29)=881 = 147

OpenStudy (anonymous):

sorry i got confuse with the problem. didnt mean too put 147

OpenStudy (skullpatrol):

Step 4 Solve the equation and find the unknowns asked for.

OpenStudy (anonymous):

ok. how do i find the unknow answer for The Pacific Paradise has______deluxe staterooms The Caribbean Paradise has____deluxe The mediterranean Paradise has___deluxe staterooms

OpenStudy (skullpatrol):

Solve the equation: p+(28+p)+(4p-29)=881 for p= the number of rooms on the Pacific

OpenStudy (anonymous):

the number of rooms on the ship is 881

OpenStudy (skullpatrol):

Solve the equation: p+(28+p)+(4p-29)=881 for p

OpenStudy (anonymous):

p=147

OpenStudy (skullpatrol):

Very good :-)

OpenStudy (anonymous):

What are the answers for all three. p= 147 The Pacific Paradise has______deluxe staterooms The Caribbean Paradise has____deluxe The mediterranean Paradise has___deluxe staterooms

OpenStudy (skullpatrol):

So now you know how many rooms are on the Pacific= 147rooms. To find the others substitute this value into: C = 28 + P M =4P - 29

OpenStudy (anonymous):

c= 28+P=147

OpenStudy (anonymous):

im still confuse a little.

OpenStudy (anonymous):

The Pacific Paradise has______deluxe staterooms The Caribbean Paradise has____deluxe The mediterranean Paradise has___deluxe staterooms

OpenStudy (skullpatrol):

Sub p=147 into c= 28+p

OpenStudy (skullpatrol):